OpenStudy (anonymous):

Simplify i^18 A. 1 B. -1 C. i D. -i

OpenStudy (anonymous):

-1

geerky42 (geerky42):

\[\Large i^1 = i \\ \Large i^2 = -1 \\ \Large i^3 = -i \\ \Large i^4 = 1\] \[\Large18/4 = 4\operatorname{r}2 ~\longrightarrow~ i^{18} = i^2 = \boxed{-1}\]